A physics instructor wants to produce a double-slit interference pattern large enough for her class to see. For the size of the room, she decides that the distance between successive bright fringes on the screen should be at least 3.00 cm. If the slits have a separation d=0.0200mm, what is the minimum distance from the slits to the screen when 632.8-nm light from a He-Ne laser is used?
Answer #1

The separation between two successive bright fringes for a double-slit interference is given by:

y=λD/d

y=λD/d

where

  • y= is the separation of two successive bright fringes in the screeny= is the separation of two successive bright fringes in the screen
  • λ=is the wavelength of the light sourceλ=is the wavelength of the light source
  • D= is the distance between the screen and the slitD= is the distance between the screen and the slit
  • d=is the slit separationd=is the slit separation


Given

  • y=3.00 cm=0.030m
  • Lamda =632.8×10−9 m
  • d=0.0200 mm=20 ×10−6 m

the slit to screen distance (D)(D) must be:

D=dy/λ

D=(20 ×10−6 m)(0.030 m)/632.8×10−9 m

D=6×10^-3/632.8×10^-9

D=0. 9481m= 94.8 cm